OpenStudy (anonymous):

x^2-8x+15=0 solve by completing the square please

OpenStudy (anonymous):

add +1 -1

OpenStudy (anonymous):

this will make the 15 a 16

OpenStudy (anonymous):

in more detail please

OpenStudy (anonymous):

x^2 - 8x + 16 -1 = 0

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

damn...ok

OpenStudy (anonymous):

usually u complete a square when u can't factorize a quadratic expression so here because we can't factorise "x^2-8x+15=0" easily we try to make the 15 a 16 and when it is 16 then we can factorize like this :....

OpenStudy (anonymous):

(x^2 - 8x + 16) -1 = 0 x^2 - 8x + 16 = 1 (x-4) (x-4) = 1 (x-4)^2 = 1 x-4 = SQRT (1) = 1 x = 1+4 = 5 makes sense? or not yet?